# Exact solution for a class of quantum models of interacting bosons

### v2

@@ II Models of interacting bosons in quantum optics
$|\alpha\rangle\equiv e^{-|\alpha|^{2}/2}\sum_{N=0}^{\infty}\frac{(\hat{a}^{\dagger})^{N}}{\sqrt{N!}}|Vac\rangle,$ (2)
@@ III Solution to the evolution problem
$\!\!\!\!|A_{3}\rangle=\left\{(\hat{A}^{\dagger})^{3}+(\hat{A}^{\dagger})^{1}\sum_{s=0}^{1}\beta_{s}\right\}\!\!|\Psi_{0}\rangle,$
$\!\!\!\!|A_{4}\rangle=\left\{(\hat{A}^{\dagger})^{4}+(\hat{A}^{\dagger})^{2}\sum_{s=0}^{2}\beta_{s}+(\hat{A}^{\dagger})^{0}\beta_{0}\sum_{s=0}^{1}\beta_{s}\right\}\!\!|\Psi_{0}\rangle,$
$\!\!\!\!|A_{5}\rangle=\left\{(\hat{A}^{\dagger})^{5}+(\hat{A}^{\dagger})^{3}\sum_{s=0}^{3}\beta_{s}+(\hat{A}^{\dagger})^{1}\sum_{s_{1}=0}^{1}\beta_{s_{1}}\sum_{s_{2}=0}^{s_{1}+1}\beta_{s_{2}}\right\}\!\!|\Psi_{0}\rangle,$
$\!\!\!\!|A_{6}\rangle=\left\{(\hat{A}^{\dagger})^{6}+(\hat{A}^{\dagger})^{4}\sum_{s=0}^{4}\beta_{s}+(\hat{A}^{\dagger})^{2}\sum_{s_{1}=0}^{2}\beta_{s_{1}}\sum_{s_{2}=0}^{s_{1}+1}\beta_{s_{2}}\right.$
$\qquad\left.+(\hat{A}^{\dagger})^{0}\beta_{0}\sum_{s_{2}=0}^{s_{1}+1}\beta_{s_{2}}\sum_{s_{3}=0}^{s_{2}+1}\beta_{s_{3}}\right\}\!\!|\Psi_{0}\rangle.$
@@
2. Reindexing the last summation to match the required form in Eq. (19) for $m+1$: To combine the remaining terms in the first summation (for $2\leq l\leq p$) with the last summation in Eq. (III), we introduce a new index $l^{\prime}\equiv l-1$ in the last summation. The new index satisfies $2\leq l^{\prime}\leq p+1=[\frac{m+1}{2}]$, which is consistent with Eq. (19). The two terms then take the form:
@@
Here, in accordance with the product over $2\leq j\leq l^{\prime}$, we have shifted the indices in $\beta$’s as follows: $s_{j}\to s^{\prime}_{j+1}$. In the second term in Eq. (III) the summation over $s^{\prime}_{2}$ has the upper limit $s^{\prime}_{1}+1\equiv s_{0}+1$, with $s_{0}=m-2l-1=(m+1)-2l^{\prime}$. This precisely matches the index in the first $\beta$-factor, which is missing in the first term of Eq. (III). Consequently, the sum over $s_{1}$ in the first term is extended to satisfy $s_{1}\leq(m+1)-2l$, as required by Eq. (19) for $m+1$. For $l=p+1$ the corresponding sum contains only $\beta_{m+1-2l^{\prime}}$, which arises from the second term in Eq. (III). Therefore, the two summations in Eq. (III) combine to reproduce the required sum over $2\leq l\leq[\frac{m+1}{2}]$, confirming that Eq. (19) holds for $m+1$.
@@
1. First Step: The first step is identical to the case of $m=2p+1$, leading to the same result as in Eq. (21). By similar arguments, this result coincides with the $l=1$ term required by Eq. (19) for $m+1=2p+1$.
2. Second Step: We now perform the same operations as in the case of $m=2p+1$, to combine the last summation in Eq. (III) with the remaining terms in the first summation (for $2\leq l\leq p$). However, since $[\frac{m-1}{2}]=p-1$, the upper limits in the two summations (after reindexing with $l^{\prime}=l-1$) coincide:
@@
$e^{-i\tau(\hat{A}+\hat{A}^{\dagger})}|\Psi_{0}\rangle=\sum_{m=0}^{\infty}\frac{(-i\tau)^{m}}{m!}\sum_{l=0}^{[\frac{m}{2}]}g^{(l)}_{m-2l}(\hat{A}^{\dagger})^{m}|\Psi_{0}\rangle$
@@ III.3 Explicit solution: the beam-splitter example
$e^{-i\tau\hat{H}_{1}}|N,0\rangle=\frac{1}{\sqrt{N!}}\left[\hat{a}^{\dagger}\cos\tau-i\hat{b}^{\dagger}\sin\tau\right]^{N}|0,0\rangle$ (36)
$=\cos^{N}\tau\sum_{n=0}^{N}\binom{N}{n}^{\frac{1}{2}}\left[-i\tan\tau\right]^{n}|N-n,n\rangle.$
The coefficients $\gamma_{n}$ appearing in the equivalent expansion of the state in Eq. (36), as given in Eq. (16) of Section II, can also be obtained directly from Eq. (36), which determines the respective quantum amplitudes as in Eq. (17). Observing that the corresponding parameter $\beta_{\ell}=(N-\ell)(\ell+1)$ leads to
@@ IV.2 Comparing the exact solution with the Gaussian squeezed state
The parametric approximation to the coefficient $\widetilde{\gamma}^{(N)}_{n}$ of the exact solution in Eq. (1), for the original quantum Hamiltonian
@@ V Concluding remarks
In this work, we have derived the exact solution for a broad class of nonlinear models describing interacting bosons, with immediate applications to the state evolution problem in quantum optics. Specifically, we have considered a pump mode propagating through a nonlinear optical medium that satisfies the phase-matching conditions, thereby generating signal mode(s) initially in the vacuum state. The solution to the state evolution problem is expressed as a series expansion in powers of the propagation time, where a single function—polynomial in the signal mode(s) populations—characterizes the specific quantum model. These results have direct applications to nonlinear models in quantum optics, including the $k$-photon down-conversion process, which plays a crucial role in advancing quantum technology.

### v4

@@ II Models of interacting bosons in quantum optics
$|\alpha\rangle\equiv e^{-|\alpha|^{2}/2}\sum_{N=0}^{\infty}\frac{(\alpha\hat{a}^{\dagger})^{N}}{N!}|Vac\rangle,$ (2)
@@ III Solution to the evolution problem
$\!\!\!\!|A_{3}\rangle=\left\{(\hat{A}^{\dagger})^{3}+(\hat{A}^{\dagger})^{1}\sum_{s_{1}=0}^{1}\beta_{s_{1}}\right\}\!\!|\Psi_{0}\rangle,$
$\!\!\!\!|A_{4}\rangle=\left\{(\hat{A}^{\dagger})^{4}+(\hat{A}^{\dagger})^{2}\sum_{s_{1}=0}^{2}\beta_{s_{1}}+(\hat{A}^{\dagger})^{0}\beta_{0}\sum_{s_{2}=0}^{1}\beta_{s_{2}}\right\}\!\!|\Psi_{0}\rangle,$
$\!\!\!\!|A_{5}\rangle=\left\{(\hat{A}^{\dagger})^{5}+(\hat{A}^{\dagger})^{3}\sum_{s_{1}=0}^{3}\beta_{s_{1}}+(\hat{A}^{\dagger})^{1}\sum_{s_{1}=0}^{1}\beta_{s_{1}}\sum_{s_{2}=0}^{s_{1}+1}\beta_{s_{2}}\right\}\!\!|\Psi_{0}\rangle,$
$\!\!\!\!|A_{6}\rangle=\left\{(\hat{A}^{\dagger})^{6}+(\hat{A}^{\dagger})^{4}\sum_{s_{1}=0}^{4}\beta_{s_{1}}+(\hat{A}^{\dagger})^{2}\sum_{s_{1}=0}^{2}\beta_{s_{1}}\sum_{s_{2}=0}^{s_{1}+1}\beta_{s_{2}}\right.$
$\qquad\left.+(\hat{A}^{\dagger})^{0}\beta_{0}\sum_{s_{2}=0}^{1}\beta_{s_{2}}\sum_{s_{3}=0}^{s_{2}+1}\beta_{s_{3}}\right\}\!\!|\Psi_{0}\rangle.$
@@
2. Reindexing the last summation to match the required form in Eq. (19) for $m+1$: To combine the remaining terms in the first summation (for $2\leq l\leq p$) with the last summation in Eq. (III), we introduce a new index $l^{\prime}\equiv l+1$ in the last summation. The new index satisfies $2\leq l^{\prime}\leq p+1=[\frac{m+1}{2}]$, which is consistent with Eq. (19). The two terms then take the form:
@@
Here, in accordance with the product over $2\leq j\leq l^{\prime}$, we have shifted the indices in $\beta$’s as follows: $s_{j}\to s^{\prime}_{j+1}$. In the second term in Eq. (III) the summation over $s^{\prime}_{2}$ has the upper limit $s^{\prime}_{1}+1\equiv s_{0}+1$, with $s_{0}=m-2l-1=(m+1)-2l^{\prime}$. This precisely matches the index in the first $\beta$-factor of the second term, which is missing in the first term of Eq. (III). Consequently, by combining the two sums, the sum over $s_{1}$ in the first term is extended to satisfy $s_{1}\leq(m+1)-2l$, as required by Eq. (19) for $m+1$. For $l=p+1$ the corresponding sum in Eq. (19) contains only $\beta_{m+1-2l}=\beta_{0}$, which arises from the second term in Eq. (III) for $l^{\prime}=p+1$. Therefore, the two summations in Eq. (III) combine to reproduce the required sum over $2\leq l\leq[\frac{m+1}{2}]$, confirming that Eq. (19) holds for $m+1$.
@@
1. First Step: The first step is identical to the case of $m=2p+1$, leading to the same result as in Eq. (21). By similar arguments, this result coincides with the $l=1$ term required by Eq. (19) for $m+1$.
2. Second Step: We now perform the same operations as in the case of $m=2p+1$, to combine the last summation in Eq. (III) with the remaining terms in the first summation (for $2\leq l\leq p$). However, since $[\frac{m-1}{2}]=p-1$, the upper limits in the two summations (after reindexing with $l^{\prime}=l+1$) coincide:
@@
$e^{-i\tau(\hat{A}+\hat{A}^{\dagger})}|\Psi_{0}\rangle=\sum_{m=0}^{\infty}\frac{(-i\tau)^{m}}{m!}\sum_{l=0}^{[\frac{m}{2}]}g^{(l)}_{m-2l}(\hat{A}^{\dagger})^{m-2l}|\Psi_{0}\rangle$
@@ III.3 Explicit solution: the beam-splitter example
$e^{-i\tau\hat{H}_{1}}|N,0\rangle=\frac{1}{\sqrt{N!}}\left[\hat{a}^{\dagger}\cos\tau-i\hat{b}^{\dagger}\sin\tau\right]^{N}|0,0\rangle$
$=\cos^{N}\tau\sum_{n=0}^{N}\binom{N}{n}^{\frac{1}{2}}\left[-i\tan\tau\right]^{n}|N-n,n\rangle.$ (36)
The coefficients $\gamma_{n}$ appearing in the equivalent expansion of the state in Eq. (III.3), as given in Eq. (16) of Section II, can also be obtained directly from Eq. (III.3), which determines the respective quantum amplitudes as in Eq. (17). Observing that the corresponding parameter $\beta_{\ell}=(N-\ell)(\ell+1)$ leads to
@@ IV.2 Comparing the exact solution with the Gaussian squeezed state
The parametric approximation to the coefficient ${\gamma}^{(N)}_{n}$ of the exact solution in Eq. (1), for the original quantum Hamiltonian
@@ V Concluding remarks
In this work, we have derived the exact solution for a broad class of nonlinear models describing interacting bosons, with immediate applications to the state evolution problem in quantum optics. Specifically, we have considered a pump mode propagating through a nonlinear optical medium that satisfies the phase-matching conditions, thereby generating signal mode(s) initially in the vacuum state. The solution to the state evolution problem is expressed as a series expansion in powers of the propagation time, where a single function – polynomial in the signal mode(s) populations – characterizes the specific quantum model. These results have direct applications to nonlinear models in quantum optics, including the $k$-photon down-conversion process, which plays a crucial role in advancing quantum technology.